OpenStudy (anonymous):

A one liter of solution of 2.5g of polystyrene in benzene has an osmotic pressure of 4.5x10^-2 atm at 50C. Calculate molar mass of polystyrene

OpenStudy (anonymous):

\[\Pi =MRT\] is what to use but i get a funky answer every time

OpenStudy (aaronq):

you have to substitute \(n=\dfrac{m}{M}\) into \(M=\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (anonymous):

Is this right? doesnt seem right \[M=\frac{ R *T }{ \Pi } \rightarrow M= 1.7x10^-3\] \[1.7x10^-3=\frac{ xmol }{ 1L } \rightarrow 1.7x10^-3 mol\] \[2.5g=\frac{ 1.7x10^-3mol }{ molar mass } = 4.25x10^-3 \]

OpenStudy (aaronq):

in the last step it should be: \(1.7*10^{-3}\;mol=\dfrac{2.5\;g}{M}\rightarrow M=\dfrac{2.5\;g}{1.7*10^{-3}\;mol}\)

OpenStudy (anonymous):

so is that 1470g/mol or 1.5x10^3 g/mol? Thank you so MUCH for your help!

OpenStudy (aaronq):

1470

OpenStudy (aaronq):

no problem dude